OpenStudy (anonymous):

Let f(x)=-4x-2 and g(x)=5x-6. find F(little black circle ) g and state it's domain

OpenStudy (anonymous):

they all are real numbers but is 6 or 1

OpenStudy (anonymous):

no domain is all real numbers. \[f(g(x))=f(5x-6)=-4(5x-6)-2\]

OpenStudy (anonymous):

now we can multiply and combine like terms \[f(g(x))=-4(5x-6)-2=-20x+24-2=-20x+22\]

OpenStudy (anonymous):

-20x

OpenStudy (anonymous):

this is a line, so domain all real numbers.

OpenStudy (anonymous):

20

OpenStudy (anonymous):

\[-20x+20\]

OpenStudy (anonymous):

34 or 12 are in the answer somewhere

OpenStudy (anonymous):

or 14

OpenStudy (anonymous):

no i don't see 34 or 14 anywhere in \[f(g(x))\]

OpenStudy (anonymous):

possible answers ..-20x^2+14x+12 all real except ..8x^2+34x-30 all real except 1 ...8x^2+34-30 all real ...-20x^2+14x+12 all real

OpenStudy (anonymous):

OOOOOOhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

\[fg(x) not f(g(x))\]

OpenStudy (anonymous):

f o g

OpenStudy (anonymous):

-20x^2+14x+12 all real except 6

OpenStudy (anonymous):

well if that is it, i am right. if it is f times g then one of those is the answer. but that open circle does not mean mulitply

OpenStudy (anonymous):

no it would not be all real except 6 because it is a polynomial so it would be all real numbers.

OpenStudy (anonymous):

no it is a close circle

OpenStudy (anonymous):

\[-20x^2+6x+8\]

OpenStudy (anonymous):

nope you are right sorry

OpenStudy (anonymous):

no you are right

OpenStudy (anonymous):

\[-20x^2+12x+12\]

OpenStudy (anonymous):

okay yeay

OpenStudy (anonymous):

domain all real numbers